OpenStudy (anonymous):

A centrifuge goes through 1550 revolutions acceleration from rest to 755 rad/s. How long did this process take? I know the answer but not the work

OpenStudy (shamim):

u hv to use 2 equations here

OpenStudy (shamim):

need equations of angular motion

OpenStudy (shamim):

bt i m giving u equations for linear motion. plz convert these into angular motion

OpenStudy (shamim):

v^2=u^2+2as v=u+at